Question 19

Suppose that a solution is prepared by mixing 25 mL (0.22 mol, 16 g) of pentane (the solute) with 45 mL (0.34 mol, 30 g) of hexane (the solvent). If the volumes add on mixing, answer questions 19-21 below.

 Calculate the molality of the pentane

a. 3.1 mol/kg

b. 4.8 mol/kg

c. 7.3 mol/kg

d. 3.1 mol/L

e. 7.3 mol/L

Question 20

Calculate the molarity of the pentane?

a. 3.1 mol/kg

b. 7.3 mol/kg

c. 3.1 mol/L

d. 7.3 mol/L

e. 45.8 mol/L

Question 21

What is the mole fraction of the pentane?

a. 0.35

b. 0.39

c. 0.64

d. 39%

e. 64 %

 Question 22

If I prepare a solution by adding 25 mL of 2.0 M NaOH into a 500 mL volumetric flask, and filling the flask to the mark with water, what will the final concentration of the solution be?

a. 0.05 M

b. 0.10 M

c. 0.5 M

d. 2.0 M

e. 40.0 M


1
Expert's answer
2022-06-07T07:32:03-0400

19.

Cmolality = mol of solute / kg of solvent

Cmolality = 0,22 mol / 0,03 kg = 7,3 mol/kg

Answer: C

20.

CM = n / V;

CM(pentane) = 0,22mol / 0,07 l = 3,1 mol/l

Answer: C


21.

Mole fraction(pentane) = n(pentane) / n(total)

Mole fraction(pentane) =0,22/(0,22+0,34)= =0,39

Answer: B


22.

n(NaOH) = CM * V=2mol/l *0,025 l =0,05mol.


for the post-dilution:

CM = n/V = 0,05mol / 0,5 l = 0,1 mol/l

Answer: B